# Lan-grasp: Using Large Language Models for Semantic Object Grasping

### v2

@@ Abstract
In this paper, we propose Lan-grasp, a novel approach towards more appropriate semantic grasping. We use foundation models to provide the robot with a deeper understanding of the objects, the right place to grasp an object, or even the parts to avoid. This allows our robot to grasp and utilize objects in a more meaningful and safe manner. We leverage the combination of a Large Language Model, a Vision Language Model, and a traditional grasp planner to generate grasps demonstrating a deeper semantic understanding of the objects. We first prompt the Large Language Model about which object part is appropriate for grasping. Next, the Vision Language Model identifies the corresponding part in the object image. Finally, we generate grasp proposals in the region proposed by the Vision Language Model. Building on foundation models provides us with a zero-shot grasp method that can handle a wide range of objects without the need for further training or fine-tuning. We evaluated our method in real-world experiments on a custom object data set. We present the results of a survey that asks the participants to choose an object part appropriate for grasping. The results show that the grasps generated by our method are consistently ranked higher by the participants than those generated by a conventional grasping planner and a recent semantic grasping approach. In addition, we propose a Visual Chain-of-Thought feedback loop to assess grasp feasibility in complex scenarios. This mechanism enables dynamic reasoning and generates alternative grasp strategies when needed, ensuring safer and more effective grasping outcomes.
@@ 1 Introduction
Objects found in household environments often require a specific way of interaction. For artificial objects, such as tools, the deployment mode can be implied by their design which ensures functionality and user safety. For instance, a knife should be typically held by the grip and not the blade. Similarly, a mug with hot tea is best held by the handle and not the rim of the mug. Incorrect handling can also impair the object itself, e.g. trying to carry a plant by the leaves would most likely lead to damage. Through experience, humans develop an intuitive understanding of objects, their parts, and their proper usage. As robots are increasingly involved in human living environments, it is crucial to provide them with the same kind of semantic knowledge.
@@
Figure 1: Robot performing the command of “Pick up the ice cream please”. The grasp on the left is generated without including semantic information while the grasp on the right is performed using our method leveraging a deeper understanding of the task and the object provided by Large Language Models.
@@
Traditional approaches to robotic grasping [1, 2, 3] only analyze the object geometry and aim to optimize the grasp stability. Without a deeper understanding of semantic aspects as described above, this can limit the usability of tools or result in object or robot impairment. Recent data-driven approaches [4, 5, 6] also account for the object class and can generate grasps appropriate for the specific object type. Several works [7, 8, 9] tackle the problem of task-specific grasping where the object is grasped differently depending on the action at hand. However, most of these methods require substantial computational resources for training and can fail to generalize to unseen object categories. Our objective is an approach for object-specific grasping that ensures tool usability and safety without any training.
@@
We proceed towards this goal by introducing Lan-grasp, a zero-shot method built on foundation models. The scale of these models and the massive size and generality of their training data allow us to reason about a large variety of objects without further training or fine-tuning. In particular, Lan-grasp uses a Large Language Model (LLM) to understand which part of an object is suitable for grasping. Next, this information is grounded in the object image by leveraging a Vision Language Model (VLM). Our method uses GPT-4 as the LLM and the OWL-ViT [10] as the VLM. However, due to the modular structure of Lan-grasp, it can easily be adapted to use other LLMs or VLMs. Finally, we use an off-the-shelf grasp proposal tool [2] to plan the grasps in accordance with the admissible parts of the object detected by the deployed foundation models [^2] Video available at https://tinyurl.com/5bnwpkuc..
@@
    We demonstrate that the presented approach can work with a wide variety of day-to-day objects without the need for additional training.
@@ 2 Related Work
Traditional grasping algorithms [1, 2, 3, 11, 12] analyze the geometry of the object and the gripper to propose and evaluate a grasping pose. Building on decades of development, these methods are fast and reliable off-the-shelf tools. However, they do not incorporate semantic information and operate based on object shape only. Also, such methods rely on a precise object model and thus suffer from partial or noisy geometry. Data-driven approaches regress grasping candidates from either single view RGB images [4, 13] or point clouds [14, 15], thus mitigating the need for a complete object model. Further, a network can learn a more natural grasping policy if human-like grasps are included in the training data, where such grasps are either created manually [16] or learned through imitation [6].
@@
Our work is closely related to task-oriented grasping (TOG) and affordance detection. TOG methods restrict the grasp candidates to a specific object part or area, conditioned on the action at hand. Muraliet al. [7] create a data set with a large number of objects and tasks and manually annotate task-specific grasp poses. Then, the authors use that data to train a grasp evaluation network. Kwaket al. [5] deploy a knowledge graph to select the gripper type and gripping force appropriate for the given object. Chenet al. [17] propose a network that jointly detects an object and generates a grasping pose according to a natural language command. However, the training requires object, grasp, and command ground truth data. Fang et al. [18] introduce TOG-Net, which optimizes task-oriented grasps and manipulation policies using simulated self-supervision.
@@
Similarly to TOG, affordance detection is the problem of identifying objects or object parts that accommodate a certain action. Doet al. [19] propose an end-to-end trained network that detects object instances in an image and assigns pixel-wise affordance masks to object parts. Liuet al. [20] build on the previous work as a backbone for affordance detection and, in addition, infer the material of object parts to further facilitate semantic grasping. Monica and Aleotti [21] propose a system that decomposes an object point cloud into meaningful parts which then serve as grasping targets. However, the part the robot has to grasp is provided by the user whereas in our method the part is suggested by an LLM. Bohg et al. [22] survey data-driven approaches to grasp synthesis, focusing on methods that sample and rank candidate grasps for both familiar and unknown objects, highlighting the role of feature extraction in these approaches. Nasiriany et al. [23] introduce a prompting framework for VLMs (PIVOT) that refines candidate actions iteratively, showing potential for spatial tasks such as grasping, but their focus is broader, addressing both navigation and manipulation tasks. Wei et al.. [24] propose a novel dataset, DexGYSNet, and utilize it to train a model for dexterous grasp generation based on language guidance. Jian et al. [25] introduce AffordPose, a large-scale dataset for affordance-driven hand-object interactions, focusing on part-level affordance labeling to guide the generation of hand-object interactions in fine detail. Zhu et al. [26] propose a framework for human-like dexterous grasping, using semantic touch codes and object functional areas to guide grasps. Ren et al. [27] introduce ATLA, a meta-learning framework that uses LLMs to accelerate tool manipulation by combining language-based policies with affordances, focusing on general tool use rather than grasping specific object parts.
@@
Foundation models have recently attracted a lot of attention in different sub-fields of robotics [28, 29, 30] and have been also applied to boost TOG and affordance detection. Ngyenet al. [31] train an open-vocabulary affordance detector for point clouds whereby CLIP is deployed to encode the affordance labels. Similarly, Tanget al. [8] use CLIP to facilitate task-specific grasping from RGB images and language instructions. The authors propose to utilize CLIP embeddings from intermediate CLIP layers to allow their affordance detector to reason about fine-grained object parts. Gaoet al. [32] annotate a large object data set with physical object properties like mass or fragility and fine-tune a VLM on it to improve manipulation planning. Other methods integrate LLMs for encoding tasks or object parts from natural language. Songet al. [33] use BERT as the language back-end and train a network that grounds object parts in a point cloud from a user instruction. Here, however, the part label is explicitly referred to in the user input. The approach of Tanget al. [9] lifts this limitation by prompting an LLM to describe the shape and parts of an object. The LLM response is then processed by a Transformer-based grasp evaluation network. Our method also relies on an LLM for deciding what object part should be grasped. The crucial difference to the above works is that our approach relies solely on foundation models and does not require any training. Thus, once more powerful foundation models are available, the performance of our approach is easily improved by switching to a novel LLM or VLM. Newbury et al. [34] conduct a systematic review of deep learning methods for six-DoF grasp synthesis, highlighting sampling-based, direct regression, and reinforcement learning methods to generate grasp poses. Wu et al. [35] propose an approach to enhance LMMs’ robustness in vision tasks, introducing reasoning capabilities to correct false premises, which improves reasoning for affordance-based grasping tasks. In a similar line of thought, Huang et al. [36] propose VoxPoser, a framework that generates 3D value maps to guide robotic manipulation using affordances extracted from LLMs and visual grounding, however their focus is on manipulation tasks rather than detecting the specific grasping part of an object.
Finally, embodied vision-language models like PALM-E [37] aim to close the gap between language, vision, and robot actions by training the network jointly on these modalities. The recent RT-2 model [38] shows remarkable capabilities of generating robot controls from language instructions. Whether these approaches choose the appropriate object parts for grasping remains an open question, since no explicit experiments were described in the publications and the code is not yet available at the time of writing.
@@ 3 Method Description
Figure 2: The core concept of our approach in a nutshell: The command from the user is turned into a prompt suitable for the Large Language Model (LLM). With this prompt as an input, the large language model outputs the proper part for grasping the object, which in this example is the cone. This word is then grounded to the object image using a Vision-Language Model (VLM). The grounded grasp part is integrated to the 3D reconstruction model of the object to generate the proper grasp.
In this section, we explain the details of our approach. Lan-grasp generates a grasping pose from an object label, a camera image showing that object, and the corresponding object geometry. The method consists of two parts. In the language module, an LLM first decides what object part to grasp, which is then grounded in the image by a VLM. The resulting bounding box is projected onto the object geometry to mark the grasp target in the grasp planning module. Thanks to the modular structure of Lan-grasp, it is easy to enhance the pipeline by leveraging more advanced models as they emerge. The pipeline of Lan-grasp is depicted in Figure 2.
@@ 3.1 Language Module
In the first step, the object label <object> provided by the user is transferred into a LLM prompt. The scheme of the prompt is chosen to be compatible with GPT-4 which is the LLM that we used in the pipeline [39]. We included the last sentence to prevent the LLM from giving extra explanations and thus only output the desired object part. We use OWL-ViT [10] as the VLM for grounding the object part label in the image. It builds on the Vision Transformer Architecture, first presented by Dosovitskiyet al. [40]. The authors then pre-train the model using contrastive learning [41] on a large image-text data set [42]. Afterward, the authors fine-tune the model on publicly available detection data sets. OWL-ViT then detects and marks the desired object part with a bounding box which is projected on the object 3D model.
@@ 3.2 Grasp Planning Module
We deploy the GraspIt! simulator [43] as our grasp proposal generator. It is a standard tool that operates on geometric models and evaluates grasps according to physical constraints. Thus, the first step for grasp planning is to create a dense 3D mesh model of the object. In our setup, we use two fixed RGB-D sensors and a turning table for object scanning. We acquire the camera poses from an Aruco board and integrate the depth images via KinectFusion [44]. However, we note that any other suitable reconstruction approach could be used here.
@@
The possible poses for grasping the object are generated by sampling. The initial gripper position is chosen based on object geometry, after that the gripper is iteratively brought closer to the object while avoiding obstacles [45]. In this regard, GraspIt! splits the scene into object and obstacle geometry, and we exploit this mechanism by marking the mesh parts that project into the VLM-generated bounding box as object and the rest as obstacle. This enforces grasping only at the desired object part. The resulting grasp proposals are ranked based on grasp efficiency and finger friction.
@@
We want to point out that our approach is agnostic about the grasp planner and could be potentially replaced by other tools that do not require a complete object model, e.g., the method of Alliegro et al. [15]. In this case, the reconstruction step could be skipped entirely and the grasp candidates could be computed on a point cloud acquired from the robot’s sensors.

### v3

@@ Abstract
In this paper, we propose Lan-grasp, a novel approach towards more appropriate semantic grasping and placing. We leverage foundation models to equip the robot with a semantic understanding of object geometry, enabling it to identify the right place to grasp, which parts to avoid, and the natural pose for placement. This is an important contribution to grasping and utilizing objects in a more meaningful and safe manner. We leverage a combination of a Large Language Model, a Vision-Language Model, and a traditional grasp planner to generate grasps that demonstrate a deeper semantic understanding of the objects. Building on foundation models provides us with a zero-shot grasp method that can handle a wide range of objects without requiring further training or fine-tuning. We also propose a method for safely putting down a grasped object. The core idea is to rotate the object upright utilizing a pretrained generative model and the reasoning capabilities of a VLM. We evaluate our method in real-world experiments on a custom object dataset and present the results of a survey that asks participants to choose an object part appropriate for grasping. The results show that the grasps generated by our method are consistently ranked higher by the participants than those generated by a conventional grasping planner and a recent semantic grasping approach. In addition, we propose a Visual Chain-of-Thought feedback loop to assess grasp feasibility in complex scenarios. This mechanism enables dynamic reasoning and generates alternative grasp strategies when needed, ensuring safer and more effective grasping outcomes.
keywords:
AI-Based Robotics, Grasping, Perception for Grasping and Manipulation
@@ 1 Introduction
Objects found in household environments often require a specific way of interaction. For artificial objects, such as tools, the deployment mode can be implied by their design, which ensures functionality and user safety. For instance, a knife should typically be held by the grip, not the blade. Similarly, a mug with hot tea is best held by the handle and not the rim of the mug. Incorrect handling can also impair the object itself, e.g., trying to carry a plant by the leaves would most likely lead to damage. Through experience, humans develop an intuitive understanding of objects, their parts, and their proper usage. As robots become increasingly integrated into human living environments, it is crucial to equip them with the same kind of semantic knowledge.
@@
Figure 1: Robot performing the command of “Pick up the ice cream please”. The grasp on the left is generated without including semantic information, while the grasp on the right is performed using our method, which leverages a deeper understanding of the task and the object provided by Large Language Models.
@@
Traditional approaches to robotic grasping (Bicchi, 1995; Miller and Allen, 2000; Ten Pas et al., 2017) only analyze the object geometry and aim to optimize the grasp stability. Without a deeper understanding of semantic aspects as described above, this can limit the usability of tools or result in object or robot impairment. Recent data-driven approaches (Jang et al., 2017; Kwak et al., 2022; Wu et al., 2023) also account for the object class and can generate grasps appropriate for the specific object type. Several works tackle the problem of task-specific grasping, where the object is grasped differently depending on the action at hand  (Murali et al., 2021; Tang et al., 2023b, a). However, most of these methods require substantial computational resources for training and can fail to generalize to unseen object categories. Our objective is to develop an approach for object-specific grasping that ensures tool usability and safety without requiring any training.
@@
We proceed towards this goal by introducing Lan-grasp, a zero-shot method built on foundation models. The scale of these models and the massive size and generality of their training data allow us to reason about a large variety of objects without further training or fine-tuning. In particular, Lan-grasp uses a Large Language Model (LLM) to understand which part of an object is suitable for grasping. Next, this information is grounded in the object image by leveraging a Vision-Language Model (VLM). Our method uses GPT-4 as the LLM and OWL-ViT (Minderer et al., 2022) as the VLM. However, due to the modular structure of Lan-grasp, it can be easily adapted to utilize other LLMs or VLMs. Finally, we use an off-the-shelf grasp proposal tool (Miller and Allen, 2000) to plan the grasps in accordance with the admissible parts of the object detected by the deployed foundation models.
We further propose an approach for putting down a grasped object in a safe manner. So far, object placing methods have been mostly focused on placement stability. However, this is not sufficient for safely handling many household items, such as cups and open bottles, which have a defined upright orientation. We argue that placing objects upright not only addresses safety issues but also tackles placement stability, as most such objects are per design stable in an upright position. We deploy the SAM 3D (Chen et al., 2025) model to generate a 3D mesh of the grasped object and retrieve the object’s pose in the input image. We then render the mesh from different views and use a VLM to select the view showing the object upright.
@@
    We demonstrate that the presented approach can work with a wide variety of day-to-day objects without requiring additional training.
@@
  5. 5.
    We propose an object placement approach that finds upright object orientations in zero-shot manner using a generative 3D foundation model and a VLM.
@@ 2 Related Work
Traditional grasping algorithms (Bicchi, 1995; Miller and Allen, 2000; Ten Pas et al., 2017; Zapata-Impata et al., 2019; Myers et al., 2015) analyze the geometry of the object and the gripper to propose and evaluate a grasping pose. Building on decades of development, these methods are fast and reliable off-the-shelf tools. However, they do not incorporate semantic information and operate based on object shape only. Also, such methods rely on a precise object model and thus suffer from partial or noisy geometry. Data-driven approaches regress grasping candidates from either single-view RGB images (Jang et al., 2017; Redmon and Angelova, 2015) or point clouds (Zhao et al., 2021; Alliegro et al., 2022), thus mitigating the need for a complete object model. Further, a network can learn a more natural grasping policy if human-like grasps are included in the training data, where such grasps are either created manually (Corona et al., 2020) or learned through imitation (Wu et al., 2023).
@@
Our work is closely related to task-oriented grasping (TOG) and affordance detection. TOG methods restrict the grasp candidates to a specific object part or area, conditioned on the action at hand. Murali et al. (2021) create a data set with a large number of objects and tasks and manually annotate task-specific grasp poses. Then, the authors use that data to train a grasp evaluation network. Kwak et al. (2022) deploy a knowledge graph to select the gripper type and gripping force appropriate for the given object. Chen et al. (2021) propose a network that jointly detects an object and generates a grasping pose according to a natural language command. However, the training requires object, grasp, and command ground truth data. Fang et al. (2020) introduce TOG-Net, which optimizes task-oriented grasps and manipulation policies using simulated self-supervision.
@@
Similar to TOG, affordance detection is the problem of identifying objects or object parts that accommodate a certain action. Do et al. (2018) propose an end-to-end trained network that detects object instances in an image and assigns pixel-wise affordance masks to object parts. Liu et al. (2020) build on the previous work as a backbone for affordance detection and, in addition, infer the material of object parts to further facilitate semantic grasping. Monica and Aleotti (2020) propose a system that decomposes an object point cloud into meaningful parts which then serve as grasping targets. However, the part the robot has to grasp is provided by the user, whereas in our method, the part is suggested by an LLM. Bohg et al. (2013) survey data-driven approaches to grasp synthesis, focusing on methods that sample and rank candidate grasps for both familiar and unknown objects, highlighting the role of feature extraction in these approaches. Nasiriany et al. (2024) introduce a prompting framework for VLMs (PIVOT) that refines candidate actions iteratively, demonstrating potential for spatial tasks such as grasping. But their focus is broader, addressing both navigation and manipulation tasks. Wei et al. (2024) propose a novel dataset, DexGYSNet, and utilize it to train a model for dexterous grasp generation based on language guidance. Jian et al. (2023) introduce AffordPose, a large-scale dataset for affordance-driven hand-object interactions, focusing on part-level affordance labeling to guide the generation of hand-object interactions in fine detail. Zhu et al. (2021) propose a framework for human-like dexterous grasping, using semantic touch codes and object functional areas to guide grasps. Ren et al. (2023) introduce ATLA, a meta-learning framework that uses LLMs to accelerate tool manipulation by combining language-based policies with affordances, focusing on general tool use rather than grasping specific object parts.
@@
Foundation models have recently attracted a lot of attention in different sub-fields of robotics (Mirjalili et al., 2023; Huang et al., 2023a; Cui et al., 2023) and have also been applied to boost TOG and affordance detection. Nguyen et al. (2023) train an open-vocabulary affordance detector for point clouds whereby CLIP is deployed to encode the affordance labels. Similarly, Tang et al. (2023b) use CLIP to facilitate task-specific grasping from RGB images and language instructions. The authors propose utilizing CLIP embeddings from intermediate CLIP layers to enable their affordance detector to reason about fine-grained object parts. Gao et al. (2024) annotate a large object data set with physical object properties like mass or fragility and fine-tune a VLM on it to improve manipulation planning. Other methods integrate LLMs for encoding tasks or object parts from natural language. Song et al. (2025) use BERT as the language back-end and train a network that grounds object parts in a point cloud from a user instruction. Here, however, the part label is explicitly referred to in the user input. The approach of Tang et al. (2023a) lifts this limitation by prompting an LLM to describe the shape and parts of an object. The LLM response is then processed by a Transformer-based grasp evaluation network. Our method also relies on an LLM to decide which object part should be grasped. The crucial difference to the above works is that our approach relies solely on foundation models and does not require any training. Thus, once more powerful foundation models are available, the performance of our approach is easily improved by switching to a new LLM or VLM. Newbury et al. (2023) conduct a systematic review of deep learning methods for six-DoF grasp synthesis, highlighting sampling-based, direct regression, and reinforcement learning methods to generate grasp poses. Wu et al. (2024) propose an approach to enhance the robustness of large multimodal models in vision tasks by introducing reasoning capabilities to correct false premises, which improves reasoning for affordance-based grasping tasks. In a similar line of thought, Huang et al. (2023b) propose VoxPoser, a framework that generates 3D value maps to guide robotic manipulation using affordances extracted from LLMs and visual grounding. However, their focus is on manipulation tasks rather than detecting the specific grasping part of an object.
Finally, embodied VLMs like PALM-E (Driess et al., 2023) aim to close the gap between language, vision, and robot actions by jointly training the network on these modalities. Vision-Language Action Models (Brohan et al., 2023; Zitkovich et al., 2023; O’Neill et al., 2024; Ghosh et al., 2024; Black et al., 2024; Kim et al., 2025) extend the idea of foundation models to robotics by training or fine-tuning large models, typically based on the Transformer architecture, on large-scale robotics datasets (O’Neill et al., 2024). They commonly employ a behavioral cloning objective to learn from human demonstrations and corresponding language instructions. A language instruction, thus, results in the execution of the corresponding skill. As the training of these models is based on demonstrations, they can be expected to implicitly align with human preferences for grasping within the domains of the training datasets. In contrast, our method explicitly aligns with human preferences, supports grasp feasibility feedback, and is not limited to a specific dataset.
The problem of object placement in robotics has been mainly approached from two directions. The first is considering the object geometry and physical properties to achieve a stable object position after placement. The second focuses on human-like object arrangement. Baumgartl et al. (2014) develop a geometry-based placement planner which finds a stable position for arbitrary objects on uneven surfaces. Haustein et al. (2019) tackle motion planning for placing household items in cluttered environments. Their method considers object stability constraints and human preferences for placing objects next to each other. More recent works integrate learning-based methods and pretrained foundation models to improve object placement. The approach of Lee et al. (2024) combines a simulation-based stability verification with a contextual placement area selection driven by an LLM. Mendez et al. (2024) deploy a diffusion model to generate a human-like object arrangement on a 2D surface and use this layout to guide object positioning. A diffusion model is also used by Nadeau et al. (2025) to generate physically stable object placements. These methods, however, do not consider the placement orientation of objects. The notion of upright or canonical object orientation has been explored in the computer vision and computer graphics communities (Fu et al., 2008; Wang et al., 2024; Lu et al., 2025). In robotics, Newbury et al. (2021) propose a two-part approach for placing an object upright. They train one CNN that regresses the angle for rotating the object upright, and one verification network that judges the current orientation. Pang et al. (2022) also propose a data-driven method for finding the upright orientation of object point clouds, where a CNN first identifies the support base plane and the upright orientation is computed from the plane normal. These models are trained on custom datasets with potential training data limitations, whereas we rely on large pretrained models like ChatGPT and SAM 3D (Chen et al., 2025).
@@ 3 Method Description
Figure 2: Our grasping approach in a nutshell: The command from the user is turned into a prompt suitable for the Large Language Model (LLM). With this prompt as an input, the LLM outputs the proper part for grasping the object, which in this example is the cone. This word is then grounded to the object image using a Vision-Language Model (VLM). The grounded grasp part is integrated into the 3D reconstruction model of the object to generate the proper grasp.
In this section, we explain the details of our method. Lan-grasp generates a grasping pose from an object label, a camera image showing that object, and the corresponding object geometry. The method consists of two parts. In the language module, an LLM first decides what object part to grasp, which is then grounded in the image by a VLM. The resulting bounding box is projected onto the object geometry to mark the grasp target in the grasp planning module. Thanks to the modular structure of Lan-grasp, it is easy to enhance the pipeline by leveraging more advanced models as they emerge. The pipeline of Lan-grasp is depicted in Fig.˜2.
@@ 3.1 Language Module
In the first step, the pipeline generates an LLM prompt from the object label <object> provided by the user. The scheme of the prompt is chosen to be compatible with GPT-4, the LLM we use in the pipeline (OpenAI, 2023). Fig.˜2 contains an example for an ice cream cone. We included the last sentence in the prompt to prevent the LLM from giving extra explanations and thus only output the desired object part. We use OWL-ViT (Minderer et al., 2022) as the VLM for grounding the object part label in the image. It builds on the Vision Transformer Architecture, first presented by Dosovitskiy et al. (2020). The authors first pre-train the model using contrastive learning (Zhai et al., 2022) on a large image-text data set (Jia et al., 2021). Afterward, they fine-tune it on publicly available object detection datasets. During inference, OWL-ViT detects and marks the desired object part with a bounding box, which we then project onto the object’s 3D model.
@@ 3.2 Grasp Planning Module
We deploy the GraspIt! simulator (Miller and Allen, 2004) as our grasp proposal generator. It is a standard tool that operates on geometric models and evaluates grasps according to physical constraints. Thus, the first step for grasp planning is to create a dense 3D mesh model of the object. In our setup, we use two fixed RGB-D sensors and a turning table for object scanning. We acquire the camera poses from an Aruco board and integrate the depth images via KinectFusion (Newcombe et al., 2011). However, we note that any other suitable reconstruction approach could be used here.
@@
The possible poses for grasping the object are generated by sampling. The initial gripper position is selected based on the object’s geometry. After that, the gripper is iteratively brought closer to the object while avoiding obstacles (Miller et al., 2003). In this regard, GraspIt! splits the scene into object and obstacle geometry, and we exploit this mechanism by marking the mesh parts that project into the VLM-generated bounding box as object and the rest as obstacle. This enforces grasping only at the desired object part. The resulting grasp proposals are ranked based on grasp efficiency and finger friction.
@@
We want to point out that our approach is agnostic about the grasp planner and could be potentially replaced by other tools that do not require a complete object model, e.g., the method of Alliegro et al. (2022). In this case, the reconstruction step could be skipped entirely, and the grasp candidates could be computed on a point cloud acquired from the robot’s sensors.
3.3 Object Placement
Figure 3: Summary of our placing approach showing one object alignment cycle. Given an RGB image of the grasped object, we first segment the object and then use SAM 3D to obtain the object 3D reconstruction and pose estimation. We then render the reconstructed mesh from six views, each showing one axis in the object frame up or down . A vision–language model is prompted to select the orientation that best matches placing the object down correctly. Finally, the robot executes the corresponding roll, pitch, and yaw rotation to align the grasp to the selected object pose.
We also address the task of determining the correct orientation of an object for placing it down. For instance, an open bottle should only be placed on its bottom part; putting it on the side or upside-down could lead to liquid spillage. Similar holds for many other household objects like mugs, glasses, plates, etc. However, ultimately, a safe placement orientation requires common-sense knowledge about each individual object. If such an orientation is meaningful for an object, we refer to it as the $up$-orientation in the following. As with grasping, we want to extract the knowledge about placement orientations from large pretrained models in a zero-shot fashion.
In our setup, we assume that the robot gripper holds an object in front of a static camera and needs to rotate the object correctly before placing it on a flat surface. The correct object orientation is determined by a VLM. Preliminary experiments have shown that generalist VLMs can detect when an object is oriented upright in the image, but struggle to reason how it should be rotated if the initial orientation is off. A brute-force solution could be to rotate the object in front of the camera in regular steps until the $up$-orientation is found. Depending on the step size, however, this procedure might take many object adjustment and VLM inference cycles until termination. Therefore, we propose a more direct approach to finding the placement orientation.
We deploy the recently released SAM 3D model (Chen et al., 2025) that generates a 3D mesh $M$ from an image and an object mask. It also estimates the transform $T_{o2c}\in SE(3)$ from the object coordinates $o$ to the camera frame. We can now render the object mesh from arbitrary view points $v\in V$, select the view $up\in V$ with the object upright via the VLM, and compute the necessary gripper rotation from $T_{o2up}T_{o2c}^{-1}$. The most general but also most inefficient strategy to construct $V$ is to sample the views uniformly from a sphere around the object. Cost-reducing heuristics could be to first compute the 3D bounding box around $M$ or the principal components of points in $M$ and assume that the object’s $up$-direction is aligned with one of the axes, thus reducing the number of views to $|V|=6$.
However, we also observe that the object frame $o$ generated by SAM 3D already fulfills this assumption, i.e., the $up$ mesh orientation is aligned with one of the axes in $o$, with the choice of the $up$-axis depending on the initial object pose in the image. We note that this is not an officially documented feature of SAM 3D, but was persistent for all objects in our experiments. We speculate that it is a byproduct of training the model on generic object mesh datasets, where the canonical mesh frame is typically upright. Further details and a comparison with the bounding box and principal component heuristics are described in Section˜4.7.
We construct $V$ such that each axis in $o$ is pointing up or down in one view with respect to the object coordination frame, i.e:
$V=\{X^{obj}_{up},X^{obj}_{down},Y^{obj}_{up},Y^{obj}_{down},Z^{obj}_{up},Z^{obj}_{down}\}$
Fig.˜3 shows the six resulting renderings. Next, these renderings are sent to the VLM to select the correct placement orientation and subsequently move the robot. We refer to one run of this pipeline as one alignment cycle. All rotations are performed in the robot world coordinate frame depicted in Fig.˜3, with the center of rotation at the gripper tool center point (TCP). The robustness of this method depends on the initial object pose estimate $T_{o2c}$. Our experiments show that SAM 3D is sufficiently reliable for estimating pitch rotations in the robot world frame, but it performs worse for roll rotations, i.e., when the object is tilted away from the camera. Therefore, after the first alignment cycle, we rotate the gripper by a $90^{\circ}$ yaw about the $z$-axis and perform a second alignment cycle to compensate for inaccuracies in the roll-angle prediction from the first cycle. After that, the object is lowered down until a contact with the table surface is detected by robot proprioception. The full alignment sequence is depicted in Fig.˜11.